A 21.34 g sample of aluminum is heated from 28 oC to its melting point of 660 oC and then melted completely. How much energy was absorbed by the sample? (Answer in kJ)

S=0.91 J/goC Delta Hf = 10.7 kJ/mol


1
Expert's answer
2021-03-23T04:21:48-0400

Solution:

Moles of Al = Mass of Al / Molar mass of Al

The molar mass of Al is 26.981 g/mol.

Hence,

Moles of Al = 21.34 g / 26.981 g mol-1 = 0.7909 mol = 0.791 mol

Moles of Al = 0.791 mol


This problem can be summarized thusly:

1) A sample of aluminum is heated from 28°C to 660°C. The heat absorbed is calculated by using the specific heat of aluminum (S = 0.91 J/g°C) and the equation: q = m × S × ΔT.

q1 = m × S × (Tf - Ti)

q1 = 21.34 g × 0.91 J/g°C × (660 - 28)°C = 12273.06 J = 12.273 kJ


2) A sample of aluminum is melted at 660°C. The heat absorbed is calculated by multiplying the moles of aluminum by the molar heat of fusion (ΔHf).

q2 = ΔHf × n = 10.7 kJ/mol × 0.791 mol = 8.464 kJ


q = q1 + q2 = 12.273 kJ + 8.464 kJ = 20.737 kJ = 20.74 kJ

q = 20.74 kJ


Answer: 20.74 kJ of energy was absorbed by the sample.